in a bag, there are coins of 25 cents, 11 cents and 5 cents in the ratio of 1:2:3. if there is $30 in all how many 5 cents coins are there?
Found 2 solutions by josgarithmetic, greenestamps:
Answer by josgarithmetic(39618)

Is "11 cents" coins a misprint? Is that really supposed to be "10 cent" coins?

Answer by greenestamps(13200)



Put the coins in groups of 1 25-cent coin, 2 10-cent coins, and 3 5-cent coins, according to the given ratio. Each of those groups of coins is worth 1(25)+2(10)+3(5) = 25+20+15 = 60 cents.

If the total value of the coins is $30 = 3000 cents, then the number of those groups of coins is 3000/60 = 50.

So then the number of 5-cent coins is 3*50 = 150.